# Is Mathematics Obsolete?

### v1

@@ 1 Mathematical Reasoning and Scientific Reasoning
Displayed prominently in the Vatican, Raphael’s fresco, The School of Athens, portrays several philosophers and thinkers of ancient Greece, arrayed around Plato and Aristotle at the center. Plato, whose philosophy of knowledge is based on forms that inhabit an extraterrestrial realm, points to the heavens. Aristotle, who took mathematical and scientific abstractions to be rooted in worldly experience, holds his hand out, open, palm facing downward. With Plato drawing our attention to the pristine beauty of mathematical forms, I imagine Aristotle saying “Whoa there, Romeo, get your head out of the clouds. We have work to do here on Planet Earth.”
@@
The scene sums up the complementary appeals of mathematical and scientific thought. Mathematics deals with abstract objects, like numbers, that are not located in time or space, whereas science is about concrete objects, located in the physical world. Mathematics is supremely rational, relying on deductive reasoning to pass from axioms and hypotheses to conclusions, whereas science is fundamentally empirical, relying on inductive reasoning to pass from data to general laws. Mathematical claims are exact, while scientific claims are approximate. Mathematical deduction yields absolute certainty, whereas scientific induction yields, at best, hypotheses that are supported by the data but can be overturned by future evidence.
These contrasts have fascinated philosophers through the ages. Historians classify the early modern philosophers of the seventeenth and eighteenth centuries as rationalists or empiricists; the rationalists took knowledge to be grounded in abstract ideas obtained by reason, independent of experience, whereas the empiricists held that the physical world is the ultimate source of all our knowledge. These characterizations evolved over the years, through the mid-twentieth century, when the logical empiricist movement classified mathematical knowledge as analytic, which is to say, true by virtue of the proper use of language and our shared linguistic framework, in contrast to the synthetic knowledge of science.
What is at stake in the philosophical debates is often murky. Sometimes, it is a matter of ontology, an account of the objects we talk about and how we should talk about them. Sometimes, it is a matter of epistemology, sorting out the appropriate ways of justifying claims to knowledge. The philosophical positions, however, often come across more as expressions of subjective preference. If you like mathematics, you are prone to view abstract mathematical knowledge as the most pristine form of knowledge and facts about the world as nothing more than pale shadows of ideal laws and mathematical truths. If you are more scientifically inclined, experience and observation are what counts, and you are more inclined to view mathematics as nothing more than the language we use to describe them.
@@ 2 Symbolic and Neural AI
Despite these ups and downs, in the latter half of the twentieth century, the symbolic worldview reigned as the dominant paradigm in our understanding of thinking and reasoning. Cognitive science was a matter of explaining intelligent human behavior by sussing out the symbolic representations and rules that give rise to it. Linguistics was a matter of working out the symbolic grammars and rules that govern language generation and interpretation. These perspectives guided efforts in mechanization. Automated reasoning meant searching for proofs in logical calculi; designing an expert system meant encoding knowledge in rules and using them to draw inferences; natural language processing meant writing grammars and parsers to analyze and generate sentences; implementing systems of computer vision meant writing algorithms to detect corners, edges, and faces, and using that data to construct symbolic representations of the underlying objects. In short, the common understanding of intelligence was that it was a matter of processing inputs into logic-based representations and operating on those. This is not to deny that the second half of the twentieth century also saw important advances in machine learning and neural networks; both were discussed at the Dartmouth workshop and were active research areas in the decades that followed. However, despite the vicissitudes of the symbolic approach, the scales always seemed to tip in its favor.
@@ 3 Is Mathematics Obsolete?
This sea change was poignant at Carnegie Mellon, where I work. In the second half of the twentieth century, Herbert Simon, a Nobel Prize winner in economics and one of the founding fathers of AI, played a tremendous role in shaping the scientific vision and direction of the university. Simon was one of the participants in the Dartmouth workshop, where he presented the Logic Theorist, a program he had written with Allen Newell and Cliff Shaw that could solve problems in propositional logic. At Carnegie Mellon, he helped found the Graduate School of Industrial Administration, the School of Computer Science, the Department of Psychology, and the Department of Philosophy. When  I arrived in 1996, logic played a prominent role in mathematics, computer science, cognitive science, philosophy, psychology, and linguistics. This was, in no small part, due to his influence. However, Carnegie Mellon was also a pioneer in machine learning. It founded the first academic department of machine learning in 2006 and offered the first undergraduate major in AI in 2018. On the university web pages, AI at CMU, you can find a history of AI at Carnegie Mellon, news and events, various degree programs, and sample curricula. Try as you may, you will not find a single occurrence of the word “logic.”
@@
That said, what distinguishes mathematics among the scholarly disciplines is the level of rigor and precision with which mathematicians state and justify their claims. Although the debate over the merits of symbolic AI vs. machine learning is not the same as the debate over the merits of mathematical vs. scientific reasoning, both are aligned with the fundamental disagreement between Plato and Aristotle as to whether finding the right concepts is a prerequisite to knowledge or whether concepts are only what emerge from empirical data. One might argue that although mathematical ideas have supported scientific exploration and practical reasoning for centuries, their importance has now diminished. Given our limited cognitive abilities, mathematical abstraction has been a valuable tool, enabling us to squeeze out additional cognitive efficiency in our attempts to navigate a complex world. But now that we have neural networks to process the data and tell us what to make of it, the idealized mathematical representations we have been using are less helpful, pale shadows of an underlying complexity that neural networks can manage more directly. One might, therefore, argue that symbolic methods are no longer important because mathematical reasoning, as we know it, is no longer important. Technology has given us something better.
Threatened with such an argument, mathematicians may retreat to aesthetics: many of us do mathematics not because it is useful, but because we enjoy it, just as we enjoy literature and art. But aesthetic and pragmatic considerations are not so easily separated. Maybe we find thinking abstractly and solving problems so enjoyable precisely because they are so generally useful; in other words, we appreciate the power of mathematics even when it is not directed at any practical goal. We might not feel the same way if mathematics were relegated to a pastime, like cricket or chess. Moreover, an appeal to aesthetics doesn’t help us solicit support for mathematics from those who do not enjoy it, and it certainly doesn’t explain why we should require our children to spend so much of their time studying it.
@@ 4 The Value of Mathematics
Generally, when we ask ChatGPT a question, we care about the answer. Whenever we turn to AI for advice, we should question the wisdom of putting our lives and livelihoods in the hands of a system we don’t understand. We should worry about whether the information we get is reliable, aligned with our interests, and likely to help us achieve our goals. We should worry about our safety and security, and since our decisions affect others, we should worry about whether the advice we get reflects our values and morals. For these reasons, we want AI to be transparent. We want AI to tell us not just what to do but also why; we want reasons, explanations, and justification. We expect no less from doctors, lawyers, salespeople, financial managers, and contractors; we ought to expect the same from AI.
@@ 5 Mathematics and AI
The good news is that the value of mathematical and symbolic reasoning has not been lost on the machine-learning community. DeepMind’s AlphaProof, using a reinforcement learning algorithm running on top of Lean, solved three of the six problems in this year’s International Mathematics Olympiad, ranking it competitive with the world’s most elite pre-college problem solvers. There has been considerable interest in the machine learning community in improving the ability of AI to carry out mathematical reasoning. There are ongoing debates between researchers as to the extent to which symbolic systems like Lean are required for a system of AI to acquire mathematical expertise, but many feel that the most promising avenues to success require a combination of neural and symbolic methods. Even ML purists recognize that obtaining mathematical arguments, explanations, and justifications is a key goal for AI, however that can be achieved.
In recent years, therefore, the phrase “AI for mathematics” has come to encompass both symbolic and machine-learning methods. My colleagues and I can breathe a sigh of relief; we can also ride the AI wave, at least for now. But we won’t always be able to ride it for free; we still need to show that symbolic methods have a vital role to play in artificial intelligence.
Let’s hope they do. Ever since Galileo proclaimed, in the sixteenth century, that the book of nature is written in the language of mathematics, mathematics has been central not only to the scientific method, but to practical decision making in several worldly pursuits, including technology, economics, finance, logistics, and public policy. Now, with the advent of AI, there are two paths we can follow. The first involves carrying out scientific reasoning and decision-making the way we have since Galileo, but using AI to do it better, improving our mathematical models and obtaining a deeper understanding of their properties. The second involves bypassing mathematics, leaving AI to draw conclusions as it sees fit, and accepting its oracular conclusions. The first path opens up exciting opportunities for mathematics and science; AI offers us new means to discover and understand phenomena that would otherwise remain opaque to us, to think and reason better, and to make better decisions. If we go down the second path, it will mean turning our back on science, relinquishing agency over our practical decisions, and giving up a vital part of what it means to be human. AI offers us the choice but does not tell us which path to take. It’s up to us to get it right.

### v2

@@ 1 Mathematical Reasoning and Scientific Reasoning
Displayed prominently in the Vatican, Raphael’s fresco, The School of Athens, portrays several philosophers and thinkers of ancient Greece, arrayed around Plato and Aristotle at the center. Plato, whose philosophy of knowledge is based on forms that inhabit an extraterrestrial realm, points to the heavens. Aristotle, who took mathematical and scientific abstractions to be rooted in worldly experience, holds his hand out, open, palm facing downward. With Plato drawing our attention to the pristine beauty of mathematical forms, I imagine Aristotle saying “Whoa there, Romeo, get your head out of the clouds. We have work to do, here on Planet Earth.”
@@
The scene sums up the complementary appeals of mathematical and scientific thought. Mathematics deals with abstract objects, like numbers, that are not located in time or space, while science is about concrete objects, located in the physical world. Mathematics is supremely rational, relying on deductive reasoning to pass from axioms and hypotheses to conclusions, whereas science is fundamentally empirical, relying on inductive reasoning to pass from data to general laws. Mathematical claims are exact, while scientific claims are approximate. Mathematical deduction yields absolute certainty, while scientific induction yields, at best, hypotheses that are supported by the data but can be overturned by future evidence.
These contrasts have fascinated philosophers through the ages. Historians classify the early modern philosophers of the seventeenth and eighteenth centuries as rationalists or empiricists; the rationalists took knowledge to be grounded in abstract ideas, obtained by reason and independent of experience, whereas the empiricists held that the physical world is the ultimate source of all our knowledge. These characterizations evolved over the years. In the mid-twentieth century, the logical empiricist movement classified mathematical knowledge as analytic, which is to say, true by virtue of the proper use of language and our shared linguistic framework, in contrast to the synthetic knowledge of science.
What is at stake in the philosophical debates is often murky. Sometimes, it is a matter of ontology, an account of the objects we talk about and how we should talk about them. Sometimes, it is a matter of epistemology, sorting out the appropriate ways of justifying claims to knowledge. The philosophical positions, however, often come across as expressions of subjective preference. If you like mathematics, you are prone to view abstract mathematical knowledge as the most pristine form of knowledge and facts about the world as nothing more than pale shadows of ideal laws and mathematical truths. If you are more scientifically inclined, experience and observation are what counts, and you are more inclined to view mathematics as nothing more than the language we use to describe them.
@@ 2 Symbolic and Neural AI
Despite these ups and downs, in the latter half of the twentieth century, the symbolic worldview reigned as the dominant paradigm in our understanding of thinking and reasoning. Cognitive science was a matter of explaining intelligent human behavior by sussing out the symbolic representations and rules that give rise to it. Linguistics was a matter of working out the symbolic grammars and rules that govern language generation and interpretation. These perspectives guided efforts in mechanization. Automated reasoning meant searching for proofs in logical calculi; designing an expert system meant encoding knowledge in rules and using them to draw inferences; natural language processing meant writing grammars and parsers to analyze and generate sentences; implementing systems of computer vision meant writing algorithms to detect corners, edges, and faces, and using that data to construct symbolic representations of the underlying objects and the relationships between them. In short, the common understanding of intelligence was that it was a matter of processing inputs into logic-based representations and operating on those. This is not to deny that the second half of the twentieth century also saw important advances in machine learning and neural networks; both were discussed at the Dartmouth workshop and were active research areas in the decades that followed. But despite the vicissitudes of the symbolic approach, the scales always seemed to tip in its favor.
@@ 3 Is Mathematics Obsolete?
This sea change was poignant at Carnegie Mellon, where I work. In the second half of the twentieth century, Herbert Simon, a Nobel Prize winner in economics and one of the founding fathers of AI, played a tremendous role in shaping the scientific vision and direction of the university. Simon was one of the participants in the Dartmouth workshop, where he presented the Logic Theorist, a program he had written with Allen Newell and Cliff Shaw that could solve problems in propositional logic. At Carnegie Mellon, he helped found the Graduate School of Industrial Administration, the School of Computer Science, the Department of Psychology, and the Department of Philosophy. When I arrived in 1996, logic played a prominent role in mathematics, computer science, cognitive science, philosophy, psychology, and linguistics, in no small part due to his influence. However, Carnegie Mellon is also a pioneer in machine learning, having founded the first academic department of machine learning in 2006 and the first undergraduate major in AI in 2018. On the university web pages, AI at CMU, you can find a history of AI at Carnegie Mellon, news and events, descriptions of various degree programs, and sample curricula. Try as you may, you will not find a single occurrence of the word “logic.”
@@
That said, what distinguishes mathematics among the scholarly disciplines is the level of rigor and precision with which mathematicians state and justify their claims. Although the debate over the merits of symbolic AI vs. machine learning is not the same as the debate over the merits of mathematical vs. scientific reasoning, both are aligned with the fundamental disagreement between Plato and Aristotle as to whether finding the right concepts is a prerequisite to knowledge or whether concepts are only what emerge from empirical data. One might argue that although mathematical ideas have supported scientific exploration and practical reasoning for centuries, their importance has now diminished. Given our limited cognitive abilities, mathematical abstraction has been a valuable tool, enabling us to squeeze out additional cognitive efficiency in our attempts to navigate a complex world. But now that we have neural networks to process the data and tell us what to make of it, the idealized mathematical representations we have been using are less helpful, pale shadows of an underlying complexity that neural networks can manage more directly. One can, therefore, argue that symbolic methods are no longer important because mathematical reasoning, as we know it, is no longer important. Technology has given us something better.
Threatened by such an argument, mathematicians may retreat to aesthetics: many of us do mathematics not because it is useful, but because we enjoy it, just as we enjoy literature and art. But aesthetic and pragmatic factors are not so easily separated. Maybe we find thinking abstractly and solving problems so enjoyable precisely because they are so generally useful; in other words, we appreciate the power of mathematics even when it is not directed at any practical goal. We might not feel the same way if mathematics were relegated to a pastime, like cricket or chess. Furthermore, an appeal to aesthetics doesn’t help us solicit support for mathematics from those who do not enjoy it, and it certainly doesn’t explain why we should require our children to spend so much of their time studying it.
@@ 4 The Value of Mathematics
Generally, when we ask ChatGPT a question, we care about the answer. Whenever we turn to AI for advice, we should question the wisdom of putting our lives and livelihoods in the hands of a system we don’t understand. We should worry about whether the information we get is reliable, aligned with our interests, and likely to help us achieve our goals. We should worry about our safety and security, and since our decisions affect others, we should worry about whether the advice we get reflects our values and morals. For these reasons, we want AI to be transparent. We want AI to tell us not just what to do but also why; we want reasons, explanations, and justification. We expect no less from doctors, lawyers, salespeople, financial managers, and contractors, and we ought to expect the same from AI.
@@ 5 Mathematics and AI
The good news is that the value of mathematical and symbolic reasoning has not been lost on the machine-learning community. DeepMind’s AlphaProof, using a reinforcement learning algorithm running on top of Lean, solved three of the six problems in this year’s International Mathematics Olympiad, ranking it competitive with the world’s most elite pre-college problem solvers. There has been considerable interest in improving the ability of AI to carry out mathematical reasoning. There are ongoing debates between researchers as to the extent to which symbolic systems like Lean are required for a system of AI to acquire mathematical expertise, but many feel that the most promising avenues to success require a combination of neural and symbolic methods. Even ML purists recognize that obtaining mathematical arguments, explanations, and justifications is a key goal for AI, however that can be achieved.
In recent years, therefore, the phrase “AI for mathematics” has come to encompass both symbolic and machine-learning methods. My colleagues and I can breathe a sigh of relief; we can also ride the AI wave, at least for now. But we won’t always be able to ride it for free; we still need to show that symbolic methods have an important role to play in artificial intelligence.
Let’s hope they do. Ever since Galileo proclaimed, in the seventeenth century, that the book of nature is written in the language of mathematics, mathematics has been central not only to the scientific method, but to practical decision making in several worldly pursuits, including technology, economics, finance, logistics, and public policy. Now, with the advent of AI, there are two paths we can follow. The first involves carrying out scientific reasoning and decision-making the way we have since Galileo but using AI to do it better, improving our mathematical models and obtaining a deeper understanding of their properties. The second involves bypassing mathematics, leaving AI to draw conclusions as it sees fit, and accepting its oracular conclusions. The first path opens up exciting opportunities for mathematics and science, because AI offers us new means to discover and understand phenomena that would otherwise remain opaque to us, to think and reason better, and to make better decisions. If we go down the second path, it will mean turning our back on science, relinquishing agency over our practical decisions, and giving up a vital part of what it means to be human. AI offers us the choice, but it does not tell us which path to take. It’s up to us to get it right.